# DesCLIP: Robust Continual Adaptation via General Attribute Descriptions for Pretrained Vision-Language Models

### v1

@@ Abstract
Continual adaptation of vision-language models (VLMs) focuses on leveraging cross-modal pretrained knowledge to incrementally adapt for expanding downstream tasks and datasets, while tackling the challenge of knowledge forgetting. Existing research often focuses on connecting visual features with specific class text in downstream tasks, overlooking the latent relationships between general and specialized knowledge. Our findings reveal that forcing models to optimize inappropriate visual-text matches exacerbates forgetting of VLMs. To tackle this issue, we propose DesCLIP, which leverages general attribute (GA) descriptions to guide the understanding of specific class objects, enabling VLMs to establish robust vision-GA-class trilateral associations rather than relying solely on vision-class connections. Specifically, we introduce a language assistant to generate concrete GA description candidates via proper request prompts. Then, an anchor-based embedding filter is designed to obtain highly relevant GA description embeddings, which are leveraged as the paired text embeddings for visual-textual instance matching, thereby tuning the visual encoder. Correspondingly, the class text embeddings are gradually calibrated to align with these shared GA description embeddings. Extensive experiments demonstrate the advancements and efficacy of our proposed method, with comprehensive empirical evaluations highlighting its superior performance compared to existing pretrained and VLM-based continual learning methods.
@@ I Introduction
In recent years, deep models pretrained on large-scale datasets have achieved remarkable success across visual, linguistic, and multi-modal domains. Pretrained vision-language models (VLMs), exemplified by CLIP [1] and ALIGN [2], have demonstrated substantial promise in handling open-vocabulary tasks. Despite their strong zero-shot capabilities in common domains, VLMs often underperform on specialized tasks, such as distinguishing low-quality images or identifying fine-grained object categories. Consequently, significant efforts [3, 4, 5, 6, 7, 8] have focused on adapting VLMs on downstream datasets to adapt to these new tasks. However, as the demand for and volume of data continue to grow, incorporating previous models and data for joint training imposes substantial storage and computational overhead. Considering the significant cost of repeatedly training foundation models, exploring continual learning (CL) becomes particularly valuable in this context.
@@
Recently, Zheng et al. [9] and Zhang et al. [10] have highlighted the risk of losing existing generalization capabilities when adjusting VLMs with generic knowledge to specialized domain models. This adjustment may result in the model losing effectiveness on prior tasks and lacking potential for optimization on subsequent tasks. This phenomenon, known as catastrophic forgetting in the field of continual learning, is particularly pronounced o n VLMs. Unlike conventional scenarios [11, 12, 13, 14], catastrophic forgetting o f VLMs impacts not only the task-specific knowledge of previously learned tasks but also the extensive pretrained knowledge, presenting significant challenges in adjusting VLMs for continual tasks.
Figure 1: (a) Existing methods: learning to match unfamiliar specific classes leads to a risk of forgetting. (b) Ours: learning to construct connections with highly relevant general attributes and gradually calibrate class-text embeddings.
Figure 2: (a) Continual fine-tuning with different task orders: familiar-first order and unfamiliar-first order. (b) Initially matching unfamiliar class texts leads to more severe forgetting, which negatively impacts the learning of subsequent tasks, resulting in overall poorer CL performance. (c) CKA [15] similarities of representation compared to a pretrained CLIP. Learning to match unfamiliar classes further disrupts the integrity of pretrained representations.
@@
Over the past few years, research [16, 10, 17, 18, 9] has explored how VLMs can adapt to CL tasks. Conventional adapting approaches [3, 4, 7, 19, 20] have shown limited effectiveness in adapting pretrained VLMs for incremental tasks. This limitation is primarily due to their reliance on shared structures and prompts across all tasks, which often results in forgetting of previously learned knowledge when accommodating new information. To address this issue, Wang et al. [16] proposed AttriCLIP, a sparse prompt selection mechanism that selects abstract attributes with high visual relevance to enrich textual hint prompts. However, these prompts depend solely on image conditions and lack association with class-relevant information, which restricts the effectiveness. Other works [21, 22, 10] have focused on selectively updating parameters of pretrained models to better support incremental continual tasks. More recently, approaches incorporating additional structures [17, 18] with sparse mechanisms have shown promise in mitigating conflicts with prior knowledge and alleviating forgetting of VLMs. Additionally, Zheng et al. [9] and Yu et al. [23] use the additional reference datasets to perform knowledge distillation, effectively mitigating the forgetting of generic knowledge.
@@
Although these studies have demonstrated some effectiveness in mitigating knowledge forgetting of VLMs, they can largely be regarded as adaptations of traditional and pretrained-vision-model-based CL methods, tailored to fit the VLM framework. Yet, they fail to fully exploit the robust cross-modal knowledge associations established during the VLM pretraining phase. For instance, these approaches overlook the visual-textual associations established via pretraining, instead relying heavily on rudimentary textual prompts (e.g., A photo of a [CLS], [XXX] [CLS]) to correlate with visual information. This introduces a risk of forgetting in unfamiliar downstream tasks. For example, CLIP may struggle with unfamiliar categories like “European garden spider,” resulting in a weak correspondence between its visual and textual representations. Forcing this association can lead to overfitting, which in turn accelerates the forgetting of pretrained and previously learned knowledge. In Fig. 2, we demonstrate that beginning with unfamiliar tasks (characterized by low image-text similarity confidence) or from familiar tasks yields markedly different evaluation results on selected subtasks of ImageNet [24]. Fig. 2 (b) underscores that forcibly aligning unfamiliar visual-text pairs hinders knowledge retention of VLMs, detracting from subsequent task learning. Additionally, Fig. 2 (c) uses centered kernel alignment (CKA) [15] to analyze representation similarity of continually tuned CLIP to a pretrained CLIP. It can be observed that matching unfamiliar classes further disrupts integrity of pretrained representation, hindering general knowledge preservation.
Another line of research [25, 26, 27, 28] have attempted to enrich textual descriptions of specific classes to assist VLMs in understanding objects in downstream tasks. However, the descriptions generated lack interaction with visual information or fail to provide class-specific representative details, making it difficult to ensure a reliable understanding of the objects.
@@
To overcome these limitations, we emphasize the transfer from generalized knowledge to specialized insights. To our knowledge, existing research has not focused on addressing forgetting by establishing robust associations between general attributes and specialized downstream class. Our approach guides the continual learning process by exploring the context encoding ability of the language branch, forming strong links between general and specialized knowledge. As shown in Fig. 1, we advocate for visual representations that align closely with highly-relevant general attribute (GA) embeddings, which are well-known to VLMs, instead of relying on naive class-text embeddings. This prevents the risk of overfitting to unfamiliar classes, as such overfitting can lead to knowledge forgetting. By gradually calibrating text embeddings to align with shared GA embeddings, we form GA-class associations for these incremental downstream tasks. In essence, we redirect the focus from conventional vision-class text connections to establishing robust vision-GA-class trilateral associations, enabling a more effective knowledge transfer that significantly mitigates forgetting of VLMs. In summary, our main contributions are as follows:
@@
  • We revisit the continual learning of VLMs, focusing on the incremental transfer from generalized to specialized knowledge. By introducing concrete descriptions of general attributes (GAs), we establish more robust vision-GA-class trilateral associations during downstream incremental phases, effectively mitigating forgetting caused by inappropriate visual-text matching.
@@
  • Our method introduces no additional overhead in terms of model structure, data replay, or feature rehearsal storage. Extensive experiments on CIFAR100 [29], ImageNet [24], and CUB-200 [30] demonstrate the exceptional performance of our approach. Thorough ablation studies and analyses further corroborate its effectiveness.
@@ II-A Continual Learning
Continual Learning (CL) investigates how deep models can incrementally learn knowledge. Existing CL research can be categorized into several types based on the strategies they employ. Among these, regularization-based methods [11, 31, 32] introduce regularization terms during model training to penalize forgetting old knowledge. These regularization terms can either focus on protecting model parameters [31] or on output distributions [11, 33] (e.g., knowledge distillation). Dynamic network-based methods [34, 35, 36, 37] aim to learn the model by introducing new structures for new tasks while preserving old knowledge, although this incurs substantial overhead as model parameters increase with the number of tasks. Recently, replay-based methods have become increasingly common. Data replay methods [38, 37] assist models in retaining old knowledge by recalling a small number of real samples. Additionally, some methods [39, 40, 41] recall old knowledge by storing sample features and the distributions of these features. However, replay-based methods introduce storage costs and require repetitive computation for old data.
In recent years, studies such as [42, 43, 44, 45] have predominantly focused on integrating additional components for incremental tasks, such as learnable prompts [44, 45, 46, 47] and adapters [42, 43], into pretrained models. This integration necessitates the development of methods for selecting and evaluating the relevance of these components to ensure both their appropriateness and compatibility with the pretrained model. However, a significant limitation arises as the number of tasks increases: the associated computational and storage costs grow substantially, posing challenges to scalability and efficiency.
@@ II-B Vision-Language Models
With advancements in pre-training techniques, large-scale foundation models [1, 48, 49, 50, 51] have significantly impacted the industry. For instance, Vision-Language Models such as Contrastive Language-Image Pretraining (CLIP) [1] and Adversarially Learned Inference for Image-Text Matching (ALIGN) [2] have demonstrated remarkable zero-shot capabilities for general tasks. However, despite being pre-trained on over 400 million image-text pairs, CLIP still face challenges in specific downstream tasks, such as accurately identifying certain types of vehicles and lizards.
@@
To better adapt VLMs for downstream tasks, various text prompt-based fine-tuning methods [3, 4, 6, 7, 52] have been proposed, which can enhance VLM performance on specific tasks. In more complex scenarios, learnable prompts can be inserted into intermediate layers [8] to incorporate more sophisticated general knowledge. Additionally, the integration of adapter structures [53, 19, 20] has also been shown to be an effective strategy. Other approaches [54, 55, 56, 57] focus on the representation alignment of VLMs and aim to improve the transfer of general knowledge. Although these methods demonstrate excellent performance in CLIP transfer tasks, they are inherently unsuitable for incrementally learning, as the additional learnable structures cannot effectively mitigate catastrophic forgetting.
II-C Continual Adaptation for VLMs
@@
Investigating the continual learning and adaptation of VLMs for diverse downstream tasks holds significant value, as it reduces data storage requirements and computational redundancy while addressing the challenge of inaccessible previous data. It is crucial to protect the model’s gerenic pretrained knowledge and previously learned knowledge. The full fine-tuning strategies discussed in II-A will lead to significant forgetting of pre-trained knowledge, which is a notable distinction between pre-trained foundation models (e.g., CLIP) and small-scale deep models. Additionally, frameworks such as CoOp [3] and CoOpOp [4] have been shown to have limited adjustment capabilities for VLMs in incremental tasks due to their reliance on shared structures and contextual prompts across all tasks, leading to forgetting old knowledge during the process of fitting new knowledge. To solve this, Wang et al. [16] introduced AttriCLIP, which establishes a shared attribute bank for all tasks and selects suitable contexts based on visual images to bridge the gap between images and text. Yu et al. [18] proposed using a mixture of experts (MoE) framework to adapt knowledge for different tasks, decoupling the model’s zero-shot capabilities from its specialized task abilities. From the perspective of parameter sparse updating, efforts from SPG [22], SparseCL [21], and SPU [10] have aimed to update VLM parameters selectively by employing appropriate “important parameter” selection patterns; for example, SPU selects more important parameters for updates based on the gradients accumulated by batches. Additionally, Zheng et al. [9] and Yu et al. [23] proposed the use of additional reference datasets to facilitate knowledge distillation in a VLM, effectively mitigating the forgetting of generic knowledge.

### v2

# DesCLIP: Robust Continual Learning via General Attribute Descriptions for VLM-Based Visual Recognition

@@ Abstract
Continual learning of vision-language models (VLMs) focuses on leveraging cross-modal pretrained knowledge to incrementally adapt to expanding downstream tasks and datasets, while tackling the challenge of knowledge forgetting. Existing research often focuses on connecting visual features with specific class text in downstream tasks, overlooking the latent relationships between general and specialized knowledge. Our findings reveal that forcing models to optimize inappropriate visual-text matches exacerbates forgetting of VLM’s recognition ability. To tackle this issue, we propose DesCLIP, which leverages general attribute (GA) descriptions to guide the understanding of specific class objects, enabling VLMs to establish robust vision-GA-class trilateral associations rather than relying solely on vision-class connections. Specifically, we introduce a language assistant to generate concrete GA description candidates via proper request prompts. Then, an anchor-based embedding filter is designed to obtain highly relevant GA description embeddings, which are leveraged as the paired text embeddings for visual-textual instance matching, thereby tuning the visual encoder. Correspondingly, the class text embeddings are gradually calibrated to align with these shared GA description embeddings. Extensive experiments demonstrate the advancements and efficacy of our proposed method, with comprehensive empirical evaluations highlighting its superior performance in VLM-based recognition compared to existing continual learning methods.
@@
Figure 1: (a) Existing methods: learning to match unfamiliar specific classes leads to a risk of forgetting. (b) Ours: learning to construct connections with highly relevant general attributes and gradually calibrate class-text embeddings, thereby significantly mitigating the forgetting caused by fitting unfamiliar class knowledge in downstream continual tasks. (c) The maintained low loss on the reference validation set demonstrates that our approach preserves the substantial potential of CL by retaining the pretrained general knowledge, thereby achieving optimal performance, compared to SL [1] and SPU [2].
@@ I Introduction
In recent years, deep models pretrained on large-scale datasets have achieved remarkable success across visual, linguistic, and multi-modal domains. Pretrained vision-language models (VLMs), exemplified by CLIP [3] and ALIGN [4], have demonstrated substantial promise in handling open-vocabulary tasks. Despite their strong zero-shot capabilities in common domains, VLMs often underperform on specialized tasks, such as distinguishing low-quality images or identifying fine-grained object categories. Consequently, significant efforts [5, 6, 7, 8, 9, 10] have focused on adapting VLMs on downstream datasets to adapt to these new tasks. However, as the demand for and volume of data continue to grow, incorporating previous models and data for joint training imposes substantial storage and computational overhead. Considering the significant cost of repeatedly training foundation models, exploring continual learning (CL) becomes particularly valuable in this context. Nowadays, CL research increasingly shifts from learning from scratch to building upon a generalist “expert” already equipped with broad knowledge, and progressively adapting it to the demands of downstream tasks. This paradigm is especially crucial for enabling continual exploration in open-world scenarios.
@@
Recently, Zheng et al. [11] and Zhang et al. [2] have highlighted the risk of losing existing generalization capabilities when adjusting VLMs with generic knowledge to specialized domain models. This adjustment may result in the model losing effectiveness on prior tasks and limiting their potential for optimization on subsequent tasks. This phenomenon, known as catastrophic forgetting in the field of continual learning, is particularly pronounced on VLMs. Unlike conventional scenarios [12, 13, 14, 15], catastrophic forgetting of VLMs impacts not only the task-specific knowledge of previously learned tasks but also the extensive pretrained knowledge, presenting significant challenges in adjusting VLMs for continual tasks.
@@
Over the past few years, research [16, 2, 17, 18, 11] has explored how VLMs can adapt to CL recognition tasks. Conventional adapting approaches [5, 6, 9, 19, 20] have shown limited effectiveness in adapting pretrained VLMs for incremental tasks. This limitation is primarily due to their reliance on shared structures and prompts across all tasks, which often results in forgetting of previously learned knowledge when accommodating new information. To address this issue, Wang et al. [16] proposed AttriCLIP, a sparse prompt selection mechanism that selects abstract attributes with high visual relevance to enrich textual hint prompts. However, these prompts depend solely on image conditions and lack association with class-relevant information, which restricts the effectiveness. Other works [21, 22, 2] have focused on selectively updating parameters of pretrained models to better support incremental continual tasks. More recently, approaches incorporating additional structures [17, 18] with sparse mechanisms have shown promise in mitigating conflicts with prior knowledge and alleviating forgetting of VLMs. Additionally, Zheng et al. [11] and Yu et al. [23] use the additional reference datasets to perform knowledge distillation, effectively mitigating the forgetting of generic knowledge.
@@
Although these studies have demonstrated some effectiveness in mitigating knowledge forgetting of VLMs, they can largely be regarded as adaptations of traditional and pretrained-vision-model-based CL methods, tailored to fit the VLM framework. Yet, they fail to fully exploit the robust cross-modal knowledge associations established during the VLM pretraining phase. For instance, these approaches overlook the visual-textual associations established via pretraining, instead relying heavily on rudimentary textual prompts (e.g., A photo of a <class name>, <prompts>+<class name>) to correlate with visual information. This introduces a risk of forgetting in unfamiliar downstream tasks (Fig. 1 (a)). For example, CLIP may struggle with unfamiliar categories like “European garden spider,” resulting in a weak correspondence between its visual and textual representations. Forcing this association can lead to overfitting, which in turn accelerates the forgetting of pretrained and previously learned knowledge. In Fig. 2, we demonstrate that beginning with unfamiliar tasks (characterized by low image-text similarity confidence) or from familiar tasks yields markedly different evaluation results on selected subtasks of ImageNet [24]. Fig. 2 (b) underscores that forcibly aligning unfamiliar visual-text pairs hinders knowledge retention of VLMs, detracting from subsequent task learning. Additionally, Fig. 2 (c) uses centered kernel alignment (CKA) [25] to analyze representation similarity of continually tuned CLIP to a pretrained CLIP. Notably, learning to match vision-language knowledge unfamiliar to CLIP consistently leads to a substantial decline in CKA at every stage, further disrupting the integrity of pretrained representations and hindering the preservation of general knowledge.
Figure 2: (a) Continual fine-tuning with different task orders: familiar-first order and unfamiliar-first order. (b) Initially matching unfamiliar class texts leads to more severe forgetting, which negatively impacts the learning of subsequent tasks, resulting in overall poorer CL performance. (c) CKA [25] similarities of representation compared to a pretrained CLIP. Task 1-5 represents ‘familiar to unfamiliar’. Learning to match unfamiliar classes further disrupts the integrity of pretrained representations.
Another line of research [26, 27, 28, 29, 30, 31] has attempted to enrich textual descriptions of specific concepts to enhance the performance of VLMs in downstream recognition tasks. However, these methods often lack explicit interaction with visual features and fail to inject fine-grained, attribute-level descriptions into the model’s optimization process in a controllable way.
@@
To overcome these limitations, we emphasize the transfer from generalized knowledge to specialized insights. To our knowledge, existing research has not focused on addressing forgetting by establishing robust associations between general attributes and specialized downstream classes. Our approach guides the continual learning process by exploring the context encoding ability of the language branch, forming strong links between general and specialized knowledge. As shown in Fig. 1, we advocate for visual representations that align closely with highly-relevant general attribute (GA) embeddings, which are well-known to VLMs, instead of relying on naive class-text embeddings. This prevents the risk of overfitting to unfamiliar classes, as such overfitting can lead to knowledge forgetting. By gradually calibrating text embeddings to align with shared GA embeddings, we form GA-class associations for these incremental downstream tasks. In essence, we redirect the focus from conventional vision-class text connections to establishing robust vision-GA-class trilateral associations, enabling a more effective knowledge transfer that significantly mitigates forgetting of VLMs in downstream recognition tasks. As shown in Fig. 1(c), our method continually supplements GA concepts to guide the VLM in adapting to new downstream tasks. This paradigm better preserves the pretrained knowledge, as evidenced by the ability to maintain a low loss on the reference validation set (i.e., ImageNet-1K[24]), thereby substantially enhancing the potential of CL scenarios. Building on this capability, we represents a novel VLM-based CL paradigm: by introducing task-agnostic general attribute knowledge, the VLM is encouraged to continually adapt to broad tasks without relying on any information from previous tasks. In summary, our main contributions are as follows:
@@
  • We revisit the continual learning of VLM-based recognition, focusing on the incremental transfer from generalized to specialized knowledge. By introducing concrete descriptions of general attributes (GAs), we establish more robust vision-GA-class trilateral associations during downstream incremental phases, effectively mitigating forgetting caused by inappropriate visual-text matching.
@@
  • Our method introduces no additional overhead in terms of model structure, data replay, or feature rehearsal storage. Extensive experiments on 6 datasets demonstrate the exceptional performance of our approach. Thorough ablation studies and analyses further corroborate its effectiveness.
@@ II-A Continual Learning
Continual Learning (CL) investigates how deep models can incrementally learn knowledge. Existing CL research can be categorized into several types based on the strategies they employ. Among these, regularization-based methods [12, 32, 33] introduce regularization terms during model training to penalize forgetting old knowledge. These regularization terms can either focus on protecting model parameters [32] or on output distributions [12, 34] (e.g., knowledge distillation). Dynamic network-based methods [35, 36] aim to learn the model by introducing new structures for new tasks while preserving old knowledge, although this incurs substantial overhead as model parameters increase with the number of tasks. Recently, replay-based methods have become increasingly common. Data replay methods [37, 36] assist models in retaining old knowledge by recalling a small number of real samples. Some methods [38, 39] recall old knowledge by storing sample features and the distributions of these features. However, replay-based methods introduce storage costs and require repetitive computation for old data. Additionally, studies [40, 41, 42] have shown that focusing on task-specific null-space projection and gradient orthogonality can lead to steady improvements in the stability–plasticity trade-off of the CL model.
In recent years, studies such as [43, 44, 45, 46] have predominantly focused on integrating additional components for incremental tasks, such as learnable prompts [45, 46, 47, 48] and adapters [43, 44], into pretrained models. This integration necessitates the development of methods for selecting and evaluating the relevance of these components to ensure both their appropriateness and compatibility with the pretrained model. To further mitigate task interference under this paradigm, Lu et al. [49] propose learning prompts in orthogonal subspaces, which helps achieve better task decoupling during continual learning. Building on this idea, Lian et al. [50] extend the subspace-based isolation strategy to the Low-Rank Adaptation (LoRA) [51] framework, a parameter-efficient fine-tuning method that introduces trainable low-rank matrices alongside frozen pretrained weights. This extension enables more effective reuse of pretrained parameters across tasks, while still preserving modularity and scalability.
@@ II-B Vision-Language Models
With advancements in pre-training techniques, large-scale foundation models [3, 52, 53, 54, 55] have significantly impacted the industry. For instance, Vision-Language Models such as Contrastive Language-Image Pretraining (CLIP) [3] and Adversarially Learned Inference for Image-Text Matching (ALIGN) [4] have demonstrated remarkable zero-shot capabilities for general tasks. However, despite being pre-trained on over 400 million image-text pairs, CLIP still faces challenges in specific downstream tasks, such as accurately identifying certain types of vehicles and lizards.
@@
To better adapt VLMs for downstream tasks, especially for visual recognition, various text prompt-based fine-tuning methods [5, 6, 8, 56] have been proposed, which can enhance VLM performance on specific tasks. In more complex scenarios, learnable prompts can be inserted into intermediate layers [10] to incorporate more sophisticated general knowledge. Moreover, text-semantic-guided learnable prompts [57] have also been shown effective in addressing domain generalization problems. Additionally, the integration of adapter structures [58, 19, 20] has also been shown to be an effective strategy. Other approaches [59, 60, 61, 62] focus on the representation alignment of VLMs and aim to improve the transfer of general knowledge. Although these methods demonstrate excellent performance in CLIP transfer tasks, they are inherently unsuitable for incremental learning, as the additional learnable structures cannot effectively mitigate catastrophic forgetting.
Recently, leveraging the fine-grained knowledge of large language models (LLMs) to enhance the vision-language alignment in VLMs has emerged as a significant research direction. This line of work [26, 27, 28, 29, 30, 31] is motivated by the observation that VLMs possess not only high-level conceptual understanding but also the capacity to interpret fine-grained visual components. Pratt et al. [26] demonstrated this capability and proposed replacing handcrafted templates with LLM-generated prompts enriched with detailed semantics. DCLIP [27] introduced multiple interpretable descriptors to provide diverse perspectives for class comprehension. Furthermore, Esfandiarpoor et al. [29] proposed FuDD, a two-stage text generation approach that uses discriminative fine-grained descriptions to reduce confusion between visually similar classes. More recently, Mirza et al. [31] developed MPVR, which employs meta-prompting strategies to produce diverse semantic texts that better support VLMs’ understanding of downstream concepts. However, while these methods focus on generating semantically rich textual inputs by leveraging the external knowledge of LLMs, they overlook the importance of controllable model optimization and pay limited attention to preserving transferable knowledge within the model itself.
II-C Continual Learning for VLMs
@@
Investigating the continual learning and adaptation of VLMs for diverse downstream tasks holds significant value, as it reduces data storage requirements and computational redundancy while addressing the challenge of inaccessible previous data. It is crucial to protect the model’s generic pretrained knowledge and previously learned knowledge. The full fine-tuning strategies discussed in II-A will lead to significant forgetting of pre-trained knowledge, which is a notable distinction between pre-trained foundation models (e.g., CLIP) and small-scale deep models. Additionally, frameworks such as CoOp [5] and CoOpOp [6] have been shown to have limited adjustment capabilities for VLMs in incremental tasks due to their reliance on shared structures and contextual prompts across all tasks, leading to forgetting old knowledge during the process of fitting new knowledge. To solve this, Wang et al. [16] introduced AttriCLIP, which establishes a shared attribute bank for all tasks and selects suitable contexts based on visual images to bridge the gap between images and text. To alleviate interference with previously acquired task knowledge in VLMs, Zhou et al. [63] proposed a novel paradigm that combines dynamic projection and fusion mechanisms, facilitating more effective utilization of cross-modal information. Furthermore, Yu et al. [18] introduced a mixture-of-experts Adapter (MoE-Adapter) framework to adaptively manage task-specific knowledge, thereby decoupling the model’s zero-shot capabilities from different specialized task adaptations. From the perspective of parameter sparse updating, efforts from SPG [22], SparseCL [21], and SPU [2] have aimed to update VLM parameters selectively by employing appropriate “important parameter” selection patterns; for example, SPU selects more important parameters for updates based on the gradients accumulated by batches. Additionally, Zheng et al. [11] and Yu et al. [23] proposed the use of additional reference datasets to facilitate knowledge distillation in a VLM, effectively mitigating the forgetting of generic knowledge.